OpenStudy (anonymous):

Solve ! y=x+3 y=-x-1

OpenStudy (anonymous):

just a real quick problem thanks for anyones help! :)

OpenStudy (alexandervonhumboldt2):

y=x+3 y=-x-1 equal them: x+3=-x-1 2x=-4 x=-2 substitute -2 for x into one of the equations y=-2+3=1

OpenStudy (anonymous):

thanks dude ! :)